OpenStudy (winsteria):

I'll give you medal and fan if you help! What is the reciprocal of –5? A. -5 B) -1/5 C) 1/5 D) 5

OpenStudy (anonymous):

B To find the reciprocal of a whole number, just turn it into a fraction in which the original number is the denominator and the numerator is 1. For example, the reciprocal of 2/3 is 3/2 (or 1-1/2) , because 2/3 x 3/2 = 1.

OpenStudy (anonymous):

@winsteria did u get it?

OpenStudy (anonymous):

reciprocal of a = 1/a so reciprocal of -5 = 1/-5

OpenStudy (winsteria):

Okay yes, and thank you both